# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v1

@@ 4.1 Optical/UV
The optical and UV light curves of AT2019aalc are shown in the bottom panel of Fig. 5. The transient was first detected by the ZTF in January with a g-band magnitude of mag. After the initial detection of the first optical flare in May (), AT2019aalc had brighten to a peak magnitude of () reached on 2019-06-18 on a timescale of roughly days. After the quickly and monotonically evolving flare, a slow decay has started, and the source was last detected with ZTF on 2021-02-18 with a magnitude of before its second flare. The re-brightening started mid-May , approximately years after the initial flare (Veres et al. 2023). The second flare has evolved very similarly to the original one. From the beginning of the second flare (the source was detected anew on 2023-05-11 at mag), a mag rise can be seen till the peak of the flare on 2023-07-13 with mag (), on a timescale of only days. Both flares evolved with constant color. However, the second flare peaked at roughly times the peak luminosity of the first flare. The peak of the second flare was followed by a short plateau phase lasted for around days. Both flares decay very slowly, on timescales of years. The first flare decayed with a power-law index of while the second has been decaying even slower with a power-law index of in -band. The decaying phases are not monotonous as bumps are tend to appear.
Prior to the discovery of AT2019aalc, the host galaxy of the transient has been monitored by several optical surveys. To study the optical variability of AT2019aalc prior to the first optical flare in , we performed aperture photometry of the All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al. 2014; Kochanek et al. 2017) using ATLAS forced photometry server (Tonry et al. 2018; Smith et al. 2020) and the Catalina Real-Time Transient Survey (CRTS; Djorgovski et al. 2011) data. In addition, Palomar Transient Factory (PTF Law et al. 2009) magnitudes were extracted from the PTF Lightcurve Table777https://irsa.ipac.caltech.edu. No significant optical variability can be seen between April and the discovery of the transient in January by ZTF (see Fig. 4).
Both optical flares of AT2019aalc detected by ZTF rising with constant color and show reddening when decaying. The second flare has similar color evolution to the first one when evolving but with a color of around the peak.
The first optical flare was not monitored in UV. The UV monitoring of the source started on 2023-06-23 around days before the optical peak of the second flare. The UV flare evolved quickly. The U-band peak is coincident with the optical peak while the shorter wavelength bands (UM2, UW1 and UW2) peaked weeks later888We note that right after the -band peak, a short plateau phase started in optical and the UV emission in the bands UM2, UW1 and UW2 peak during this plateau. The transient is most luminous at the shortest wavelength ranges covered by UVOT. The UV emission peaks around erg/s in filter UW2. After the peak epoch, the UW2U color shows reddening which suggests increasing absorption or a drop of the blackbody temperature. Towards the end of the UVOT monitoring, the measured UM2 magnitudes decrease back to the level of the baseline GALEX-NUV magnitudes measured between and .
@@ 4.2 Infrared
We detected AT2019aalc in all of the three bands observed with the WIRC. The observed magnitudes are plotted in the top panel of Fig. 5. The host galaxy of AT2019aalc is presented in the 2MASS All-Sky Catalog of Point Sources (Cutri et al. 2003) as 2MASS 152416660451191. The observed magnitudes in April are higher with mag in J-, and Ks-bands and with mag in H-band indicating the clear brightening of the IR counterpart with respect to archival data.
The top panel of Fig. 5 also shows the long-term light curve starting just before the first optical flare. There are clear signs of variability throughout the infrared light curve starting in February with the distinct first dust echo flare that peaks around six months after the first optical peak. The rise of the dust echo of the second flare becomes visible in the latest WISE data release. The light curve varies by around mag in W1 and mag in W2 before the dust echo flare, which is consistent with statistical AGN variability (Berk et al. 2004). In contrast, the dust echo presents an increase of mag in W1 and mag in W2. The first detection of the dust echo of the second flare shows a brightening of mag in both bands and can be expected to increase further in upcoming data releases. Because the brightness is significantly above the limiting magnitude where Eddington bias becomes an issue for stacked single-exposure photometry (Necker et al. 2024), we can estimate the pre-flare mid-IR color using the median of the light curve prior to the dust echo flare with . Although there is clear AGN activity prior to the dust echo flare, this is still below the widely used AGN identification cut of (Stern et al. 2012). However, according to the reliability of WISE selection as a function of a simple W1W2 color selection (Stern et al. 2012), the color of AT2019aalc suggests a reliability of only %. With the onset of the dust echo, the IR emission becomes dominated by the heated dust, which is shown in Fig. 6. The change from bluer to redder colors come from cooling of the dust. The re-brightening of the second dust echo is consequently accompanied by another color change towards the blue.
@@ 4.3 X-ray
AT2019aalc was monitored with the Swift/XRT at keV energies for weeks with weekly cadence around the peak of the optical flare and later for additional months with weekly and bi-weekly cadences. The luminosity evolution is shown in the bottom panel of Fig. 5 while the count rate variability is plotted in the top panel of Fig. 7. The first two visits show an increase of the X-ray emission. This flaring episode peaked at on 2023-06-29 (MJD) (see Fig. 5.) before it experienced a very quick decreasing. This flare peaked weeks before the optical peak. The count rate dropped by a factor of , from to , within a timescale of days. Later, the X-ray emission shows limited variability for several months, however, less significant flaring episodes appear that peaked in the end of September (MJD) and mid-February (MJD). The latter was followed by a bump in the optical and UV light curves around weeks later while no optical/UV data is available around the another X-ray increasing. In the end of April , a rapid and extreme flaring episode was detected in X-rays that lasted less than a month but reached a luminosity of erg/s which is almost a magnitude larger compared to the peak of the first X-ray flaring episode. Interestingly, this X-ray flare was also accompanied by an increasing phase in the optical weeks later, according to the most recent ZTF observations999https://lasair-ztf.lsst.ac.uk/objects/ZTF19aaejtoy/. Notably, days passed between both the first and second and the third and fourth X-ray flaring episodes. While no data is available for a longer period between the second and third flares when the source was too close to the Sun.
In Fig. 8 we show the source spectrum derived from the full observing campaign. The spectrum is dominated by a power-law component but a blackbody component is also significantly presented. The spectral index of the stacked spectrum fitted with a power-law+blackbody model is while the blackbody temperature is eV. These imply an overall unabsorbed flux of F erg cm-2s-1 corresponding to a luminosity value of L erg/s. The count rate of the observations is cts s-1. The photon index variability of AT2019aalc is presented in the bottom panel of Fig. 7.
AT2019aalc (= SRGe J152416.7045118) was observed by the eROSITA (Sunyaev et al. 2021) soft X-ray telescope four times starting from 2020-02-02 (roughly half year after the peak of the first optical flare) with a months cadence. The X-ray light curve reached a plateau between August and January with a flux of F erg cm-2s-1 in the energy range of keV. The source had a soft thermal spectrum described with a blackbody temperature of eV (van Velzen et al. 2024). One observation with XMM-Newton Observatory (Jansen et al. 2001) has been performed in 2021-02-21 during a period of optical quiescence. The best-fit power-law index and the observed keV flux of this observation were and F erg cm-2s-1 (68 confidence), respectively (Pasham 2023). These imply that the two X-ray flares observed with the Swift/XRT peaked at much higher luminosities than the plateau or long-term flaring episode between the optical flares, however, there is no X-ray data from around the peak of the first optical flare. Furthermore, the eROSITA results confirm the presence of a blackbody component in the X-ray spectrum, which is very soft and dominating the variability of the X-ray source.
@@ 4.4.1 Temporal evolution
The host galaxy of AT2019aalc was detected in the NRAO VLA Sky Survey (NVSS, White et al. 1997) in February 1995 and in the FIRST survey in February 2000 (both at 1.4 GHz) with peak flux densities of mJy and mJy respectively, suggesting no substantial variability of the host galaxy prior to the discovery of the transient event.
Later, the transient’s sky location has been observed during the first and second epochs of the Very Large Array Sky Survey (VLASS). The VLASS covers the entire sky visible to the Karl G. Jansky Very Large Array (VLA) at GHz wide band offering an angular resolution of arcsec and an rms of mJy beam-1 in each epoch (Lacy et al. 2020). One of the main goals of the VLASS multi-epoch survey is to detect various types of extragalactic radio transients, such as TDEs, accretion state changing AGNs and core-collapse supernova (Lacy et al. 2020; Zhang et al. 2022). To date, the first two epochs of the survey have been fully published. Our source has peak flux density values of mJy/beam (on 2019-03-14) and mJy/beam (on 2021-11-06), corresponding to 3 GHz radio luminosities of and in the Epoch 1 and Epoch 2 images, respectively. This factor of increase in flux from the first VLASS epoch (three months before the optical peak) to the second VLASS epoch (2 years post-peak) is statistically significant (at the 8-level, as estimated from the rms in the VLASS “Quick Look” images, van Velzen et al. 2024).
In addition, we found a radio source at the position of AT2019aalc in the first data release (DR1) of the Rapid ASKAP Continuum Survey (RACS; McConnell et al. 2020). The RACS DR1 includes initial observations made for RACS-low (central frequency of MHz)101010https://www.cadc-ccda.hia-iha.nrc-cnrc.gc.ca/en/racs/. Moreover, the source has been observed in two different fields and, consequently, at two different epochs. The observed peak flux densities of mJy/beam (on 2019-04-24) and mJy/beam (on 2020-04-30) correspond to a MHz radio luminosity of and indicate that the radio flare seen in the VLASS data started at least days after the peak of the first optical flare, sometime between April 2020 and November 2021.
Interestingly, the radio luminosity in the beginning of our ATCA monitoring is at the same level within uncertainties ( ) when the source was observed for the second time in the VLASS in April 2021. Since the radio brightness does not show rapid variability during the months of our ATCA monitoring, we can reasonably assume that the sharp increasing resulted in a long-lasting plateau. Our multi-frequency monitoring with the ATCA implies no substantial variability on monthly timescales at GHz while it shows slow decreasing at GHz and at GHz over time. We observed the radio source at GHz at one epoch. We present the ATCA radio multi-frequency light curves of AT2019aalc together with the archival observations in the middle panel of Fig. 5 and focusing on the ATCA monitoring in Fig. 9.
@@ 4.4.2 The radio spectrum
To further characterize the radio emission of AT2019aalc, we derived the radio spectral index of the radio source in each epoch of our ATCA monitoring. The spectral index between GHz and GHz (, defined as ) does not show any obvious signs of variability on a timescale of months. We give an average spectral index of . We calculated a pre-flare spectral index using the VLASS epoch and RACS epoch flux densities (these flux values are explained in Sect. 4.4.1.). The spectral index we derived between GHz and GHz this way is .
Our high-frequency K-band observations centered at GHz, however, do not follow the steep power-law and suggest a turnover between GHz and GHz. The spectrum of high frequency peaked radio sources are usually explained with significant absorption taken place in the system due to synchrotron self-absorption (although free-free absorption may also play a role, (e.g., Berton et al. 2020). We further discuss the features of the radio spectrum in Subsect. 5.3.
@@ 4.4.3 VLBI detection
The final naturally-weighted VLBI map of AT2019aalc is shown in Fig. 11. We detected a single radio-emitting feature with a signal-to-noise ratio (SNR) of with our GHz EVN+e-MERLIN VLBI observation. We obtained the following parameters of the fitted component; a total flux density of mJy and a size of mas at a position angle of . We calculated the uncertainties of the flux density and size parameters following the formulas of Fomalont (1999). The minimum resolvable size () of a Gaussian component fitted to naturally weighted VLBI data was calculated following Kovalev et al. (2005). This yields mas implying that the fitted Gaussian component is resolved. We note that at high SNR, can be significantly smaller than the size of the resolving beam (Bertero & de Mol 1996; Kovalev et al. 2005) due to applying a specific a priori hypothesis about the shape of the emitting region which is in our case the two-dimensional Gaussian used to fit the observed brightness distribution of the radio source.
The positional uncertainty was estimated as follows. When calculating the VLBI positional uncertainty, we took into account both the positional uncertainty of the phase calibrator source and its angular separation from our target source. The position of the phase calibrator source is listed with an error of mas in the third realization of the International Celestial Reference Frame (ICRF3; Charlot et al. 2020). The angular separation between the target and the calibrator source signifies a positional error of about mas at GHz (Chatterjee et al. 2004). The resolution of the observation depending on the thermal noise of the interferometric images (Reid et al. 1988) has been also considered. Following Reid et al. (1988), we estimated a thermal noise-limited positional accuracy of mas. The resultant positional uncertainty of mas is therefore dominated by the uncertainty arising from the angular separation between the calibrator and the target source. The radio position agrees with the source position in the Gaia DR3 (Gaia Collaboration et al. 2023) catalog within the positional uncertainties.
@@
where is the flux density of the fitted Gaussian component measured in Jy, and are the major and minor axes full width at half maximum (FWHM) of the fitted component in mas, and is the observing frequency in GHz. The obtained brightness temperature is K at GHz. The brightness temperature is an effective parameter that is commonly used in radio astronomy to describe the physical properties of emitting material in astrophysical objects (e.g., Lobanov 2015). It can be determined by imaging and modeling the structure of a given emitting region. The derived value is crucial to characterize the detected radio emission, which can be either thermal or non-thermal.
@@ 4.5 Spectroscopic results
We present the archival and the newly obtained optical spectra of AT2019aalc in Fig. 12 with marking the notable emission lines. In the following, we highlight the remarkable line detection of the host spectrum observed before the first optical flare (pre-flare spectrum), the one observed between the two flares (post-flare-1 spectrum) and spectra that were taken during the second flare (over-flare-2 spectra).
@@ 4.5.1 Pre-flare spectrum
A spectrum of the host galaxy of AT2019aalc was taken in April by the SDSS spectrograph. We present this spectrum with the best fits for the stellar and gas components in Fig. 13.
We identify strong Balmer emission lines. These lines are generally thought to be originating from the broad-line region i.e. the vicinity of the SMBH. We estimate the FWHM of the Balmer lines to FWHM(Balmer) kms-1 for the broad component of the Balmer lines (see Fig. 13 subplots) in the host spectrum, which is typical for an unobscured, broad line AGN. The line widths (FWHM(Balmer) km s-1) presented in the SDSS DR7 broad-line AGN catalog (Liu et al. 2019) for our source are consistent with this classification. In addition to the Balmer lines, we detect the recombination line He i which is also produced in the BLR.
The remarkable [O iii] and [O iii] forbidden lines exhibited already in the SDSS spectrum indicate that the host AGN was not completely dormant before the optical flares, which is further supported by the radio detections prior to the first flare (Subsect. 4.4.1). Other forbidden emission lines commonly seen in AGN such as [S ii] , [S ii] and [N ii] are detected as well. These lines are produced in the narrow-line region (NLR) in AGN. We note that the sulfur lines are also present in the spectra of the TDE AT2019dsg, which was hosted by a quiescent galaxy. We detect the [O ii] [O ii] line doublet, which is a potential indicator of star formation or AGN-driven outflows (e.g., Santoro et al. 2020).
@@ 4.5.2 Post-flare-1 spectrum
The first LRIS spectrum was taken almost years after the first flare and years before the second one.
We detected an increase in the Balmer lines series and the forbidden lines [O iii] and [O iii] . The Balmer lines are also remarkably broadened.
In addition to the strong Balmer emission lines, we can identify the He ii He ii and transitions and the Bowen Fluorescence line O iii . Wavelength coincidences between emission lines (”line fluorescence”) can be an important source of radiative excitation. In the case of Bowen Fluorescence, the He ii–O iii and N iii coincidences lead the mechanism. The BF lines occur due to the excitation in certain states of O iii and N iii by the absorption of He ii Lyman photons. These excited states lead to a cascade of transitions that can be observed as emission lines in the optical and NUV regimes (Netzer 1990; Trakhtenbrot et al. 2019). Interestingly, the existence of BF emission lines in astronomical objects has been predicted by Netzer et al. (1985) as the required gas densities are consistent with what is expected from the extremely dense environment in some AGN and gaseous nebulae.
We detect the high ionization coronal lines [Fe vii], [Fe x] and [Fe xiv]. The very high ionozation potentials of eV and eV of [Fe xiv] and [Fe x] , respectively, point to very energetic processes at work. First Oke & Sargent (1968) studied coronal lines in detail in a Seyfert galaxy (NGC 4151). These lines are commonly detected in AGN and commonly thought to be powered by photoionization of the NLR (Cerqueira-Campos et al. 2021). In some extreme cases, a shock model i.e. collosional ionization for the coronal line region is required to explain the origin of these lines as it is the case in the Seyfert galaxy NGC 1068 (Oliva 1997).
We identify the [Ne V] emission lines at and in the spectrum. These lines are not strongly excited in the inner regions of AGNs, but are prominent emission lines in the lower density extended emission line regions (EELRs) excited by the AGN via photoionization or shocks. The forbidden line [O iii] also appears, however, bending from H is possible.

### v2

@@ 4.1 Optical/UV
The optical and UV light curves of AT2019aalc are shown in the bottom panel of Fig. 3. The transient was first detected by the ZTF in January with an -band magnitude of mag. AT2019aalc brightened to a peak magnitude of () reached on 2019-06-18 on a timescale of days. The second flare started mid-May , approximately years after the initial flare (Veres et al. 2023). The second flare evolved very similarly to the original one. From the beginning of the second flare, a mag rise was seen up to the peak of the flare on 2023-07-13 with mag (), on a timescale of days. The second flare peaked at roughly times the peak luminosity of the first flare. The peak of the second flare was followed by a short plateau phase that lasted for around days. Both flares declined very slowly, on timescales of years. The first flare decayed with a power-law index of while the second with in the -band. The decaying phases are not monotonic, as bumps tend to appear. During the second optical flare, these bumps appear four times: end of December , mid-March , end of May and mid-August (see Figure 3). Both optical flares brightened with constant color () and showed reddening when decaying.
Prior to the discovery of AT2019aalc, the host galaxy of the transient was monitored by several optical surveys. To study the pre-discovery optical variability of AT2019aalc, we performed aperture photometry of the All-Sky Automated Survey for Supernovae (ASAS-SN: Shappee et al. 2014; Kochanek et al. 2017) and the Catalina Real-Time Transient Survey (CRTS: Djorgovski et al. 2011). In addition, Palomar Transient Factory (PTF Law et al. 2009) magnitudes were extracted from the PTF Lightcurve Table666https://irsa.ipac.caltech.edu. No significant optical variability is visible between April and the discovery of the transient in January by ZTF (see Fig. 4).
The UV monitoring of the source started on 2023-06-23 around days before the optical peak of the second flare. The U-band peak is coincident with the optical peak while the shorter wavelength bands peaked weeks later. We note that right after the -band peak, a short plateau phase started in optical and UV. The transient is most luminous at the shortest wavelength ranges covered by UVOT. The UV emission peaks around erg/s in the filter UW2. After the peak, the UW2U color shows reddening. Towards the end of the UVOT monitoring, the measured UM2 magnitudes decreased back to the level of the baseline GALEX-NUV magnitudes measured between and .
@@ 4.2 Infrared
We detected AT2019aalc in all of the three bands observed with the WIRC. The observed magnitudes are plotted in the top panel of Fig. 3. We compare these to pre-flare magnitudes presented in the 2MASS All-Sky Catalog of Point Sources (Cutri et al. 2003). The observed magnitudes in April are higher with mag in J-, and Ks-bands and with mag in H-band indicating the clear brightening of the IR counterpart with respect to archival data.
The top panel of Fig. 3 shows the light curve starting just before the first optical flare. The distinct first dust echo flare that peaks around six months after the first optical peak. The rise of the dust echo of the second flare becomes also visible in the latest WISE data releases. The light curve varies by around mag in W1 and mag in W2 before the dust echo flare, which is consistent with statistical AGN variability (Berk et al. 2004). In contrast, the dust echo presents an increase of mag in W1 and mag in W2 after the first optical flare. Following the second optical flare, the IR light curve shows a brightening of around mag in both bands and is increasing when the optical flare is already decaying. Because the brightness is significantly above the limiting magnitude where Eddington bias becomes an issue for stacked single-exposure photometry (Necker et al. 2024), we can estimate the pre-flare mid-IR color using the median of the light curve prior to the dust echo flare with . Although there is clear AGN activity prior to the dust echo flare, this is still below the widely used AGN identification cut of (Stern et al. 2012). However, according to the reliability of WISE selection as a function of a simple W1W2 color selection (Stern et al. 2012), the color of AT2019aalc suggests a reliability of only %. With the onset of the dust echo, the IR emission becomes dominated by the heated dust. The change from bluer to redder colors comes from cooling of the dust. The re-brightening of the second dust echo is consequently accompanied by another color change towards the blue shown in Fig.5.
@@ 4.3 X-ray
The X-ray luminosity evolution is shown in the bottom panel of Fig. 3 while the count rate variability is plotted in the top panel of Fig. 6. The first flaring episode peaked at in the end of June (see Fig. 3.) This flare peaked weeks before the optical peak. The count rate dropped by a factor of within a timescale of days. Later, the X-ray emission shows limited variability for several months. A hint of increased emission was detected in mid-February which was also followed by a bump in the optical and UV light curves around weeks later. In the end of April , a rapid and extreme flaring episode started in X-rays that lasted less than a month but reached a luminosity of erg/s which is almost a magnitude larger compared to the peak of the first X-ray flaring episode. Interestingly, this X-ray flare was also accompanied by a bump in the optical weeks later. A similarly luminous and rapid flare was detected in August and once more, accompanied by an optical bump weeks later. Around days passed between the second and third and the third and fourth X-ray peaks.
In Fig. 7 we show the source spectrum derived from the observing campaign. The spectral index of the stacked spectrum fitted with a power-lawblackbody model is while the blackbody temperature is eV. These imply an overall unabsorbed luminosity value of L erg/s. The mean count rate of the observations is cts s-1. The photon index variability of AT2019aalc is presented in the bottom panel of Fig. 6.
AT2019aalc (= SRGe J152416.7045118) was observed by the eROSITA (Sunyaev et al. 2021) soft X-ray telescope four times starting from 2020-02-02 (roughly half a year after the peak of the first optical flare) with a months cadence. The X-ray light curve reached a plateau between August and January with a flux of F erg cm-2s-1 in the energy range of keV. The source had a soft thermal spectrum described with a blackbody temperature of eV (van Velzen et al. 2024). One observation with XMM-Newton Observatory (Jansen et al. 2001) was performed on 2021-02-21 during a period of optical quiescence. The best-fit power-law index and the observed keV flux of this observation were and F erg cm-2s-1 (68 confidence), respectively (Pasham 2023).
@@ 4.4.1 Temporal evolution
The pre-flare system was detected in the NRAO VLA Sky Survey (NVSS, White et al. 1997) in February 1995 and in the FIRST survey in February 2000 (both at GHz) with peak flux densities of mJy and mJy respectively, suggesting no substantial variability. These flux densities correspond to a monochromatic radio luminosity of W Hz-1 measured at GHz which is typical for Seyfert galaxies (Meurs & Wilson 1984).
Later, the transient’s sky location was observed in the Karl G. Jansky Very Large Array Sky Survey (VLASS, Lacy et al. 2020). The VLASS observes the entire sky visible to the VLA at GHz wide band three times. Our source was detected with peak flux density values of mJy/beam (on 2019-03-14), mJy/beam (on 2021-11-06) and mJy/beam (on 2024-07-20), corresponding to 3 GHz radio luminosities () of , and , respectively. This factor of increase in flux from the first VLASS epoch (three months before the first optical peak) to the second VLASS epoch (2 years post-peak) is statistically significant (at the 8-level, as estimated from the rms in the VLASS “Quick Look” images, van Velzen et al. 2024).
In addition, we found a radio source at the position of AT2019aalc in the first data release of the Rapid ASKAP Continuum Survey (RACS: McConnell et al. 2020). The source was observed in two different fields and, consequently, at two different epochs. The observed peak flux densities of mJy/beam (on 2019-04-24) and mJy/beam (on 2020-04-30) correspond to a MHz radio luminosity of and indicate that the radio flare seen in the VLASS data started at least days after the peak of the first optical flare, sometime between April 2020 and November 2021.
The radio luminosity in the beginning of our ATCA monitoring was at the same level ( ) when the source was observed for the second time in the VLASS in April 2021. Since the radio brightness does not show rapid variability during the months of our ATCA monitoring, we can reasonably assume that the sharp increasing resulted in a long-lasting plateau. We present the ATCA radio multi-frequency light curves of AT2019aalc together with the archival observations in the middle panel of Fig. 3.
@@ 4.4.2 The radio spectrum
To further characterize the radio emission of AT2019aalc, we derived the radio spectral index of the radio source in each epoch of our ATCA monitoring. The spectral index between GHz and GHz (, defined as ) does not show any obvious signs of variability on a timescale of months. We give an average spectral index of . We calculated a pre-flare spectral index using the VLASS epoch and RACS epoch flux densities . The spectral index we derived between GHz and GHz this way is .
Our high-frequency K-band observations centered at GHz, however, do not follow the steep power-law and suggest a turnover between GHz and GHz.
@@ 4.4.3 VLBI detection
The naturally-weighted VLBI map of AT2019aalc is presented in Fig. 9. We detected a single radio-emitting feature with a signal-to-noise ratio (SNR) of . We obtained the following parameters of the fitted component: a total flux density of mJy and a size of mas at a position angle of . We calculated the uncertainties of the flux density and size parameters following the formulas of Fomalont (1999). The minimum resolvable size () of a Gaussian component fitted to naturally weighted VLBI data was calculated following Kovalev et al. (2005). This yields mas implying that the fitted Gaussian component is resolved.
@@
where is the flux density of the fitted Gaussian component measured in Jy, and are the major and minor axes full width at half maximum (FWHM) of the fitted component in mas, and is the observing frequency in GHz. The obtained brightness temperature is K at GHz. The brightness temperature is an effective parameter that is commonly used in radio astronomy to describe the physical properties of emitting material in astrophysical objects (e.g., Lobanov 2015).
@@ 4.5 Spectroscopic results
We present the archival and the newly obtained optical spectra of AT2019aalc in Fig. 17 with marking the notable emission lines. In the following, we highlight the remarkable line detection of the pre-flare spectrum observed before the first optical flare, the one observed between the two flares (post-flare-1 spectrum) and five spectra that were taken during the second flare (over-flare-2 spectra).
@@ 4.5.1 Pre-flare spectrum
A spectrum of the pre-flare system was taken in April by the SDSS spectrograph. We present this spectrum with the best fits for the stellar and gas components in Fig. 18.
We identify strong Balmer emission lines. We calculated FWHM(Balmer) kms-1 for the broad component of the Balmer lines in the pre-flare spectrum, which is typical for an unobscured, broad-line AGN. The line widths (FWHM(Balmer) km s-1) presented in the SDSS DR7 broad-line AGN catalog (Liu et al. 2019) for our source are consistent with this classification.
Several forbidden lines such as [O iii] and [O iii] , [S ii] , [S ii] and [N ii] were detected. These lines are produced in the narrow-line region (NLR) in AGN. We detected the [O ii] [O ii] line doublet, which is a potential indicator of AGN-driven outflows (e.g., Santoro et al. 2020).
The pre-flare spectrum also shows the detections of two high-ionization coronal lines ([Fe vii] and [Fe xiv]) which are further discussed below.
@@ 4.5.2 Post-flare-1 spectrum
The first LRIS spectrum was taken almost years after the first optical flare and years before the second one.
In this spectrum, prominent Balmer lines were detected, similarly to the SDSS spectrum. The spectrum exhibits the aforementioned forbidden lines detected in the archival spectrum.
We identify the Bowen Fluorescence line O iii and He ii .
The high ionization coronal lines [Fe vii], [Fe x], [Fe xiv] and [Fe vii] were detected. We identify the [Ne V] emission lines at and in the spectrum. Note that the pre-flare SDSS spectrum does not cover these wavelengths, therefore these lines are not necessarily related to the optical flares discussed here. These lines are prominent emission lines originating from the inner NLR excited by the AGN via photoionization or shocks and are sometimes considered as coronal lines (e.g., Cann et al. 2020).